OpenStudy (anonymous):

FAN+5MEDALS Ami bought some gardening tools. She paid $15 for a rake, $18 for a garden hose, and $47 for a wheelbarrow. What percent of the total amount that Ami paid was the price of the rake? Round to the nearest percent, if necessary. %

OpenStudy (anonymous):

58%

OpenStudy (anonymous):

15+18+47 = 80 15/80= 0,1875 18.75 = 19%

OpenStudy (anonymous):

@WinterRaine We have the same name but i dont have the Winter part. And the only reason im randomly telling you this is because i have not seen anyone with my name, ever.

OpenStudy (anonymous):

dude its for rake not wheelbarrow

OpenStudy (anonymous):

15+18+47 = 80 15/80 = your answer. Just in decimal form.

OpenStudy (anonymous):

WinterRaine and Reptile_Girl777 THATS not how you do, teach correctly please

OpenStudy (anonymous):

Reptile_Girl777 and Reptile_Girl777 Did 100% wrong, for those who give them medal, please learn math before you comment